n a​ state's Pick 3 lottery​ game, you pay ​$0.87 to select a sequence of three digits​ (from 0 to​ 9), such as 411. If you select the same sequence of three digits that are​ drawn, you win and collect ​$297.12. Complete parts​ (a) through​ (e). a. How many different selections are​ possible? 1000 b. What is the probability of​ winning? . 001 ​(Type an integer or a​ decimal.)

c. If you​ win, what is your net​ profit? ​$ nothing ​(Type an integer or a​ decimal.)
Answer #1

(a) Number of different selections = 103 = 1000

(b)

Probability of winning= 1/1000 =0.001

(c)

Event Value (x) Probability (p) x.p
Win 297.12 - 0.87 = 296.25 0.001 296.25 X 0.001 = 0.29625
Loss - 0.87 0.999 - 0.87 X 0.999 = - 0.86913
Total 0.29625 - 0.86913 = - 0.57288 (Loss)

If I win, my net profit = 0.29625

So,

Answer is:

0.29625